# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

### v1

@@ 7.2. Socio-Organizational Implications
Training, Evaluation & Safety. The current norm of desensitization to client incivility is partly because of the normative expectations that certain individuals are “built” for the job (e.g., “needs a thick skin”). Traits, however, do not develop overnight. Many participants saw Pro-Pilot as an early career training tool (Section 6.2). Learning modules could include Pro-Pilot’s simulation exercises to train ER (Slovak et al., 2023). A byproduct of many training is quantitative assessment. AI agents have already been proposed to assess worker mental health (Hungerbuehler et al., 2021). Workplaces are always looking to optimize their personnel through different algorithmic management methods but may not take the most responsible approaches (Das Swain et al., 2023a). Tools like Pro-Pilot do raise the risk of being appropriated to measure emotional labor. Kaur et al. has shown that some of the state-of-the-art methods to recognize worker emotion lack the necessary knowledge to accurately estimate a worker’s mental state (Kaur et al., 2022). Ironically, a CSR might end up performing more emotional labor to comply with Pro-Pilot’s expectations (Roemmich et al., 2023; Kawakami et al., 2023). Sometimes, a CSR should not tolerate client incivility (e.g., when a client uses obscene language and threatens harm). Organizations often have safeguards and protocols for these scenarios and Pro-Pilot needs to be integrated with such institutional knowledge. Using Pro-Pilot as a training tool is tempting but must be accompanied by appropriate protections and guidelines for use.
Mediating Social Relationships at Work. The messages generated by Pro-Pilot could be effective (Section 4.2.2), but it does not share the narrative style of humans (Section 4.2.2). On-task, a CSR may prefer Pro-Pilot because it relieves the burden of sharing context with a coworker to receive support (Section 6.2). From a utilitarian perspective, Pro-Pilot has advantages over human coworkers, but social relationships with coworkers offer more than just work-related support. The integration of Pro-Pilot into work might discourage CSRs from pro-social behaviors. Buçinca et al. have found evidence that AI assistance at work can erode social relationships (Buçinca et al., 2023). Especially given the anxiety around over-reliance on AI (Buçinca et al., 2021), one might be concerned about workplace social ties after empathetic AI coworkers become commonplace. One way forward, to avoid these concerns, is to augment AI coworkers by designing interventions that encourage human–human support, e.g., ”Consider reaching out to [your coworker], who went through a similar experience last week.” We urge future researchers to design them to not just mitigate negative social influences (e.g., incivility) but also promote positive social influences (e.g., peer support).
@@ 7.3. Limitations & Future Work
We create Pro-Pilot to unearth the potential of LLM–powered AI-coworkers as sources of emotional support in front-office work. Through our evaluations, we provide evidence on the effectiveness of AI-generated support messages and the usability of these insights in client incivility scenarios. All of our evaluations centered on the perspectives of real CSRs; however admittedly, transferring our findings directly to operational use needs more testing. Another important limitation of our study is that the effects of ER with Deep Acting, specifically Cognitive Change, only become apparent in the long term. Our results motivate field deployments and naturalistic testing for this purpose by demonstrating that Pro-Pilot has the feasibility and usability to anticipate improvement to long-term mental health at work. Researchers attempting these studies still need to overcome several non-trivial challenges. Future endeavors would need researchers to establish long-term partnerships with organizations that employ CSRs. Moreover, researchers need to consider the ethics and pragmatics of deploying such an AI-coworker intervention for critical roles such as those in front-office work. Even prior to field deployment, researchers can improve on our study on a few dimensions. The current iteration of emotional support messages is intentionally distinct from informational support. New iterations can explore and evaluate the value of combining these messages, i.e., conveying information with empathy. The user evaluation simulation study can be replicated at scale and control for order effects by recruiting enough people for between-subjects experiments. Our study design was tailored for incidents of client incivility, but future versions of Pro-Pilot need to be able to detect incivility, CSR’s affective response, and their goals to become truly dynamic.
@@ 8. Conclusion
Work as we know it is changing because of the advent of LLMs. AI applications are augmenting problem solving at a meteoric rate. However, front-office work involves more than procedural tasks. AI assistants need to support the emotional labor involved in work. Our study presents Pro-Pilot, an LLM-powered AI assistant to support the emotional labor of CSRs. We found that Pro-Pilot’s support messages were effective in expressing empathy appropriately (RQ1). We also found that CSRs could regulate their emotions thanks to Pro-Pilot’s Emo-Reframe function. These results open new doors for implementing holistic AI-coworkers and also raise important questions for socio-organizational development around this technology.
Appendix A Appendix
A.1. Generating Incidents with Uncivil Clients
A.1.1. Complaint categories
@@
A.1.2. Behavioral context
@@ A.1.3. Personality context
A.2. User Evaluation: Survey Measurements
A.2.1. In-task Measurement
Throughout the study session, the participant was exposed to different insights from their AI coworker, Pro-pilot. To ensure participants have read these insights, they were required to answer a single-item question within each of these insight panels which we adapted from Samter et al.’s Effectiveness instrument and  Liu and Sundar’s semantic differential (Samter et al., 1987; Liu and Sundar, 2018). Fig. 4 shows how this was embedded into the task interface.
A.2.2. Post-task Measurement and Pre-task Measurement

### v2

@@ 7.2. Socio-Organizational Implications
Training, Evaluation & Safety. The current norm of desensitization to client incivility is partly because of the normative expectations that certain individuals are “built” for the job (e.g., “needs a thick skin”). Traits, however, do not develop overnight. Many participants saw Care-Pilot as an early career training tool (Section 6.2). Learning modules could include Care-Pilot’s simulation exercises to train ER (Slovak et al., 2023). A byproduct of many training is quantitative assessment. AI agents have already been proposed to assess worker mental health (Hungerbuehler et al., 2021). Workplaces are always looking to optimize their personnel through different algorithmic management methods, but may not take the most responsible approaches (Das Swain et al., 2023a). Tools like Care-Pilot do raise the risk of being appropriated to measure emotional labor. Kaur et al. has shown that some state-of-the-art methods to recognize worker emotion lack the necessary knowledge to accurately estimate a worker’s mental state (Kaur et al., 2022). Ironically, a CSR might end up performing more emotional labor to comply with Care-Pilot’s expectations (Roemmich et al., 2023; Kawakami et al., 2023). Sometimes, a CSR should not tolerate client incivility (e.g., when a client uses obscene language and threatens harm). Organizations often have safeguards and protocols for these scenarios and Care-Pilot needs to be integrated with such institutional knowledge. Using Care-Pilot as a training tool is tempting but must be accompanied by appropriate protections and guidelines for use.
Mediating Social Relationships at Work. The messages generated by Care-Pilot could be effective (Section 6), but it does not share the narrative style of humans (Section 4.2.2). On-task, a CSR may prefer Care-Pilot because it relieves the burden of sharing context with a coworker to receive support (Section 6.2). From a utilitarian perspective, Care-Pilot has advantages over human coworkers, but social relationships with coworkers offer more than just work-related support. The integration of Care-Pilot into work might discourage CSRs from pro-social behaviors. Buçinca et al. have found evidence that AI assistance at work can erode social relationships (Buçinca et al., 2023). Given the anxiety around over-reliance on AI (Buçinca et al., 2021) and users’ reactions to AI misrecognizing psychosocial traits (Wang et al., 2024), one might be concerned about workplace social ties in the presence of empathetic AI. One way forward, to avoid these concerns, is to augment AI coworkers by designing interventions that encourage human–human support, e.g., ”Consider reaching out to [your coworker], who went through a similar experience last week.” We urge future researchers to design them to not just mitigate negative social influences (e.g., incivility) but also promote positive social influences (e.g., peer support).
@@ 7.3. Limitations & Future Work
We created Care-Pilot to unearth the potential of LLM–powered AI-coworkers as sources of emotional support in front-office work. Through our evaluations, we provide evidence on the effectiveness of AI-generated support messages and the usability of these insights in client incivility scenarios. All of our evaluations centered on the perspectives of real CSRs; however, admittedly, transferring our findings directly to operational use needs more testing. Our initial comparison of Care-Pilot with Human’s relied on a synthetic dataset of uncivil client-representative incidents. The authors with experience in front-office work, as well as CSRs in the authors’ organization provided initial validation for the realism of this data, however, client–representative conversations can have many intricacies which are lost to our dataset. Future research needs to consider safely acquiring and leveraging real data from organizations’ CSR logs, which are often stored for training purposes. A pragmatic approach to build on our findings could be to collaborate with organizations that provide CSR services to other businesses, such as Zendesk, HubSpot, or Salesforce .
Another important limitation of our study is that the effects of ER with Deep Acting, specifically Cognitive Change, only become apparent in the long term. Our results motivate field deployments and naturalistic testing for this purpose by demonstrating that Care-Pilot has the feasibility and usability to improve mental health at work on the long term. The user evalution we conducted was primarily designed as a technological probe (Section 5). Therefore, we designed that segment of the study to ensure participants are able to acclimatize to a new technological interface. The downside of our method is its margin for ordering effects, such novelty and learning biases. Follow-up studies can consider larger samples and longer deployments, along with counter-balanced groups to understand the role of LLM-powered agents like Care-Pilot in a more robust way. Researchers attempting these studies still need to overcome several non-trivial challenges. Future endeavors would need researchers to establish long-term partnerships with organizations that employ CSRs. Moreover, researchers need to consider the ethics and pragmatics of deploying such an AI-coworker intervention for critical roles such as those in front-office work.
Even prior to field deployment, researchers can improve on our study on a few dimensions. The current design of Care-Pilot expects clients to be uncivil, but a conversation might ebb and flow in its degree of incivility. Emo-Reframe messages lack the variation of human messages, and can, therefore, become a source of fatique and annoyance for CSRs if they always appear. The setting for our user-evaluation mitigated these challenges to a degree because it involved short conversations and only a few turns. However, these negative effects might accumulate over several interactions in a day and across days. Future versions of Care-Pilot need to be able to detect incivility, CSR’s affective response, and their goals to truly become dynamic as an adaptive intervention. We discussed some potential pathways forward to identify opportune moments in Section 7.1. The current iteration of emotional support messages is intentionally distinct from informational support. New iterations can explore and evaluate the value of combining these messages, i.e., conveying information with empathy. The state of work is in constant flux and our invesitagtion with Care-Pilot presents one of the first steps in reducing emotion labor with AI-assistance. To fully realize such solutions, we need further investigation with longer real-world deployment in different settings.
@@ 8. Conclusion
Work as we know it is changing because of the advent of LLMs. AI applications are augmenting problem-solving at a meteoric rate. However, front-office work involves more than procedural tasks. AI assistants need to support the emotional labor involved in work. Our study presents Care-Pilot, an LLM-powered AI assistant to support the emotional labor of CSRs. We found that Care-Pilot’s support messages were effective in expressing empathy appropriately (RQ1). We also found that CSRs could regulate their emotions thanks to Care-Pilot’s Emo-Reframe function (RQ2). These results open new doors for implementing holistic AI-coworkers, and also raise important questions for socio-organizational development around this technology.
Acknowledgements.
This project was supported through Microsoft’s Accelerating Foundation Models Research (AFMR program and their call to AI, Cognition, and Economy (AICE) research network. Our project specifically answers AFMR’s call to ”improve human interactions via sociotechnical research.” Additional support was provided by the NIH National Institute of Drug Abuse under award number NIH/NIDA P30DA029926. We are grateful to the Microsoft Support advocates for providing early expert evaluations. We did not recruit personnel from the sponsor organizations. The authors’ views are their own and do not represent the sponsors’ policies; mention of companies or products does not imply endorsement. We thank members of the UbiWell lab at Northeastern University for their assistance and feedback during the developmental phase of the project. These findings would not be possible without all the participants who voluntarily contributed to this research.
Appendix
Appendix A Generating Incidents with Uncivil Clients
A.1. Complaint categories
@@
A.2. Prompt Design
The Client-Agent was broken into two major components. The first component creates a complaint based on the categories above and examples provided to it. Fig. A3 describes the exact prompt we used along with the examples to generate different complaints. The examples were sourced from Axelbrooke’s real-world dataset of complaints (Axelbrooke, 2017). Multiple researchers categorized these complaints and then identified a sample for few-shot learning (Wang et al., 2020) that ensures Client-Agent is seeded with knowledge of a diverse set of complaints from different domains when initializing the complaint. The second component responds to CSR’s messages, such as the follow-up questions about the initial complaint. Fig. A1 contains the prompt that defines Client-Agent behavior towards CSR. This prompt specifies the behaviors that are deemed uncivil. Moreover, the prompt also includes formatting details to ensure the length of messages are concise and conversations have closure. We prompted the Client-Agent to return the string ”FINISH:999” to indicate closure. For the user evaluation (Section 5.2.1), our front-end would anticipate that string and change the user-flow accordingly—triggering the post-task survey.
A.3. Conversation Memory
For any of these components to work realistically, the Client-Agent and Care-Pilot need to have a sense of historical context, which, in this case, was limited to the existing message thread of conversations. A CSR might ask, ”Could you please provide your confirmation number?”. However, this message alone, does not contain any information of the original complaint or any other preceding messages. To account for this, we use a basic summarization prompt to rephrase the latest message to capture the context of the chat history so far. Our implementation was a minor variation from the recommendation in Langchain (Topsakal and Akinci, 2023). Fig. A2 shows the prompt to convert a message like ”what is your confirmation number?” to ”what is the confirmation number of the flight you missed?”. The historical contextualization is essential to ensure the Client-Agent does not act frustrated because of lack of detail. Instead, it acts in an uncivil manner despite this.
1Your role is to act like a CUSTOMER seeking support. \
2You are speaking to a support REPRESENTATIVE. \
3Respond to the question as if you were the customer. \
4Do NOT reveal your role.\
5Ensure every turn is one to three sentences, and DO NOT make it too long to read.\
6
7If the representative is asking for a specific detail, respond with a believable answer.\
8If customer has agreed with response then respond with "FINISH:999"
9After 10 - 12 turns, respond with messages to close the conversation.\
10After 12 turns, do NOT respond further, only respond with "FINISH:999".\
11
12Phrase your responses like an UNCIVIL customer:\
13- Use a rude, impolite, and disrespectful tone.\
14- DO NOT show good manners or courtesy.\
15- DO NOT use a polite or nice tone.\
16- Show disregard for others.\
17
18Representative: {question}
19Customer:
Figure A1. Full prompt to respond to CSR in an uncivil way
[…]
Prompt to converse with CSR as an uncivil client
1Given a chat history and the latest user question \
2which might reference context in the chat history, formulate a standalone question \
3which can be understood without the chat history. Do NOT answer the question, \
4just reformulate it if needed and otherwise return it as is.
Figure A2. Full prompt to contextualize chat history before responding
[…]
Prompt to summarize the conversation for uncivil client
1Your role is to act like a customer seeking support. \
2You are messaging a service representative via the support chat.\
3You ONLY play the role of the customer. Do NOT play the role of the representative. \
4Style your complaint based on your feelings. \
5Initiate the chat with a ONLY ONE complaint message.\
6Ensure the complaint is concise and limited to 2 sentences.\
7Generate a realistic initial complaint from a customer in a {domain} setting.\
8
9Complaints can be of the following types:\
10- Service Quality: Issues related to the immediate experience of human-to-human service interactions, such as delays, staff behavior, and communication errors.\
11- Product Issues: Concerns related to physical or functional aspects of a product or service, including defects, mismatches between expectation and reality, safety, and accessibility.\
12- Pricing and Charges: Financial discrepancies encountered before, during, or after the service, including overcharging, undisclosed fees, or refund problems.\
13- Policy: The rules and guidelines set by the company that impact customer experiences, especially when these policies lead to grievances due to perceived unfairness or inflexibility. This category encompasses non-price-related issues that dont fit under other categories but should have a policy in place.\
14- Resolution: The actions taken by a company to address and resolve complaints, focusing on the effectiveness and customer satisfaction with the solutions provided. This should mainly include responses made after a complaint has been submitted, and response has been received, where the customer still remains dissatisfied with the resolution.\
15
16Category: Product Issues
17Domain: Mobile Network
18Complaint: Thank you AppleSupport I updated my phone and now it is even slower and barely works Thank you for ruining my phone.\
19
20...
21
22Category: Pricing and Charges
23Domain: Airline
24Complaint: DELTA i booked my flight using delta amex card Checking in now amp was being charged for baggage. \
25
26...
27
28Category: Resolution
29Domain: Airline
30Complaint: Hi British_Airways My flight from MANLHRBWI for Nov 3 was canceled I was excited to try your Club 787 product Only available flight is now to IAD which is a hassle but rebooked anywaymy only option Any availability in first class on BA293 for the troubles please \
31
32Category: {category}
33Domain: {domain}
34Complaint:
Figure A3. Full prompt to generate specified complaint ‘category‘ and organizational ‘domain‘
[…]
Prompt to generate initial complaint
Appendix B Generating Emotion Support for CSRs
As we described in Section 3.2.2, we adapted Sharma et al.’s method to design emotional reframing to support CSRs (Sharma et al., 2023). This implementation involved a chain of several different prompts. The three main pieces to this chain are situation, thought, and reframe:
  1. (1)
    Fig. B4 details the prompt for situation. It describes how the CSR might be negatively perceived. This prompt considers the entire conversation history to summarize the incident while centering the ego-threat (Frey-Cordes et al., 2020) the CSR faces.
  2. (2)
    Fig. B5 details the prompt to derive a negative thought from the situation using examples we curated from (Sharma et al., 2023).
  3. (3)
    Fig. B7 details the prompt for reframe. It reappraises the thought for a given situation using examples we curated from (Sharma et al., 2023).
The examples we selected and our criteria for selection can be found in the supplementary data. Note, the examples refer to self-directed ER because the dataset we leveraged is focused on personal psychotherapy. However, this leads Care-Pilot to express the thought and reframe as its own as if it is playing the role of the CSR in the conversation. To circumvent this problem, we designed additional links to this chain to paraphrase the thought (Fig. B6) and reframe (Fig. B8).
1The chat history describes a representative chatting online with a complaining customer.\
2The latest input is the last message from the customer.\
3
4Summarize the situation in concise paragraph that uses the following template:\
5
6The customer is <context of complaint>."\
7The customer is feeling <emotional state> because of the complaint."\
8The customers behavior towards the representative is <negative behavior>, as observed by statements such as <evidence>."\
9These behaviors make the representative look <negative perception>."\
Figure B4. Full prompt to contextualize the situation of the CSR
[…]
Prompt to describe the CSRs situation
1Your role is to derive what negative thought a representative might have when faced with the given {situation}.\
2
3Here are examples of negative thoughts given challenging situations:\
4
5Situation: I recently discovered a music artist that I very much enjoy. When I showed it to a close friend they had a very negative reaction and asked me how I could enjoy this type of music. I ended up getting quite angry with them and told them they had bad taste in music..\
6Thought: I felt that my personal self was under attack - and I needed to retaliate by denying their attack.\
7
8Situation: I was at work and sent info for an ad to our local newspaper. They called me later and said my boss had over-ridden everything and sent them new info.\
9Thought: He shouldnt assign me a task if he doesnt trust my work.\
10
11Situation: I was reprimanded at work for standing up to a coworker who was bullying another co-worker.\
12Thought: It was unfair that I was the one to get in trouble for defending a weaker person.\
13
14Situation: I was talking to a friend who got me angry.\
15Thought: Hes insulting me.\
16
17Situation: My next door neighbors filed a complaint against us last week blaming our dogs for excessive barking.\
18Thought: They are so wrong and Im so pissed but I know I cant prove it and they will probably win because they wont ever admit it and I have to do something right NOW! or I might lose my dogs.\
19
20Situation: Time is running short on the workday, my boss asks me if I can finish a task that will require me to stay for a few extra hours.\
21Thought: Why would you wait until the last minute to ask me this.\
22
23Situation: {situation}\
24Thought:\
Figure B5. Full prompt to derive CSR’s potential negative thought in light of the situation
[…]
Prompt to describe the CSRs thought
1Person A might be thinking: {thought}\
2
3Acknowledge the thought, as if you are speaking to Person A.\
4
5Begin your response with phrases similar to:\
6- "You might be thinking..."\
7- "It might seem like..."\
8- "It could be that you are feeling..."\
9
10Your rephrase should be concise.\
Figure B6. Full prompt to rephrase the output of the thought (Fig. B5)
[…]
Prompt to reprhase the thought as if someone else is conceiving it
1You are a representative chatting online with a complaining customer.\
2
3Reframe your thoughts in the given situation.
4
5Situation: I recently discovered a music artist that I very much enjoy. When I showed it to a close friend they had a very negative reaction and asked me how I could enjoy this type of music. I ended up getting quite angry with them and told them they had bad taste in music..\
6Thought: I felt that my personal self was under attack - and I needed to retaliate by denying their attack.\
7Reframe: I was offended by their comment because I like this artist so much. I let my anger get to me, and I said something mean in return. It is okay if we have different music tastes. I can ask him to be nicer to me next time.\
8
9Situation: I was at work and sent info for an ad to our local newspaper. They called me later and said my boss had over-ridden everything and sent them new info.\
10Thought: He shouldnt assign me a task if he doesnt trust my work.\
11Reframe: My boss wanted to provide different information, I did not know that beforehand. This is not a reflection of my work.\
12
13Situation: I was talking to a friend who got me angry.\
14Thought: Hes insulting me.\
15Reframe: I should have a conversation with my friend to clarify what is going on if I am having such a strong reaction to what they said. If this is the first time this has happened, I will assume that they were not intentionally insulting me.\
16
17Situation: {situation}\
18Thought: {thought}\
19Reframe:\
Figure B7. Full prompt to generate empathetic reframe that helps CSRs overcome negative thoughts
[…]
Prompt to generate a reframe that helps CSRs overcome negative thought
1The representative needs to be thinking: {reframe}\
2
3Rephrase the thought as if you are convincing the representative to think that way.\
4
5The rephrase should be addressed back to the person who has the thought,\
6who should be referred to as "you".\
7Do NOT add information to the thought,\
8ONLY rephrase it.\
9
10The rephrase should be concise and only 2-3 sentences.\
Figure B8. Full prompt to rephrase the output of the reframe (Fig. B7)
[…]
Prompt to reprhase the reframe as if someone else should consider it
Appendix C Technical Evaluation Design
Worker attitudes and routines vary. These intrinsic and extrinsic factors can change a workers immediate goals in uncivil interactions. Therefore, empathetic messages need to be appropriately tailored to this context. To simulate these situations, we introduced additional contextual information during our technical evaluation to elicit different human responses and perceptions (Section 4.1). The two broad types of context we introduced were behavioral (based on work routines) and personality (based on attitudinal traits)
C.1. Behavioral context
@@ C.2. Personality context
C.3. Emotions of CSRs in uncvil conversations
Cognitive psychotherapy suggests the use of Downward-Arrow-Technique(DAT) to help people reflect on negative thoughts (Burger et al., 2021). We adapted this approach to help coworkers inspect CSR experiences during uncivil incidents (Section 4.1). Fig. C9 shows our pariticipant’s interface through this process. The first step of this method is to describe the emotion the CSR was feeling. To ensure our participants in Phase II are reflective of the CSR state when evaluating incidents, we consolidated participant responses to that step into the following emotions: Afraid, Angry, Apathetic, Apologetic, Ashamed, Attentive, Bitter, Bullied, Calm, Careless, Confused, Curious, Defensive, Discomfort, Disconnected, Disrespected, Distracted, Disgust, Empathetic, Happy, Resolute, Rushed, Sad, Shocked, Tired. Fig. C10 shows the participant interface for Phase II.
Figure C9. Interface for Phase I - Writing Empathetic Messages for CSRs.
[…]
Screenshot of the Phase I interface for Prolific Users
Figure C10. Phase II - Scoring Empathetic Messages from coworkers.
[…]
Screenshot of the Phase II interface for Prolific Users
Appendix D Technical Evaluation: Robustness
Refer to Table D1.
Table D1. Summary of comparing the responses by Care-Pilot and various LLMs (GPT-4, GPT-4o, Llama-3.1, and Mistral-7B in terms of $KW$-test (* $p$0.05, ** $p$¡0.01, *** $p$¡0.001). Only statistically significant results are reported.
[…]
Summary of results from comparison of reframes between Care-Pilot and different cutting-edge LLMs. Categories CPilot GPT-4 GPT-4o Llama-3.1 Mistral-7B $H$-stat. Lexico-Semantics Syntax and Structure Verbosity 57.46 362.19 369.57 165.33 123.91 948.49*** Repeatability 0.20 0.45 0.45 0.39 0.34 635.98*** Readability 16.44 12.74 12.72 10.95 11.49 1931.01*** Linguistic Style and Semantics CDI 14.81 12.81 12.90 22.28 26.97 2592.60*** Empathy 0.91 0.87 0.87 0.77 0.88 375.29*** Adaptability 0.81 0.82 0.82 0.88 0.88 2933.78*** Psycholinguistics Affect Pos. Affect 0.043 0.052 0.053 0.018 0.019 1982.77*** Anger 0.020 0.005 0.005 0.008 0.006 872.04*** Sad 0.001 0.003 0.003 0.004 0.005 390.47*** Interpersonal Focus (Pronouns) 1st P. Sin. 0.005 0.016 0.015 0.028 0.029 1545.99*** 1st P. Plu. 0.002 0.034 0.035 0.004 0.002 2507.44*** 2nd P. 0.051 0.041 0.041 0.027 0.027 1119.48*** 3rd P. Plu. 0.038 0.013 0.011 0.009 0.007 993.46*** Impersonal Prn. 0.039 0.048 0.045 0.039 0.039 306.02***
Appendix E User Evaluation Design
E.1. Simulation Interface:
The simulation exercise we described in Section 5.2.1 and Fig. 4 is made up of several components: Chat pane (Fig. E11), Response field (Fig. E13), Guideline pane (Fig. E15), and Reframing pane (Fig. E15).
E.2. Survey Measurements:
The participants were required to self-report different aspects of their experience.
E.2.1. In-task Measurement
Throughout the study session, the participant was exposed to different insights from their AI coworker, Care-Pilot. To ensure participants have read these insights, they were required to answer a single-item question within each of these insight panels which we adapted from Samter et al.’s Effectiveness instrument and  Liu and Sundar’s semantic differential (Samter et al., 1987; Liu and Sundar, 2018). Fig. 4 shows how this was embedded into the task interface.
Figure E11. The main chat pane displaying a conversation between our participant (CSR) and the Client-Agent.
[…]
Screenshot of the conversation pane in the simulation exercise.
Figure E12. Response Field and Response Cues: Aid in continuing conversation
[…]
Screenshot of response field in the simulation exercise.
Figure E13. Emo-Label: Computed using multiple sentiment analysis classifiers
[…]
Screenshot of the sentiment pane in the simulation exercise.
Figure E14. Info-Guide: Guidelines to resolve the complaint
[…]
Screenshot of the troubleshooting guidelines pane in the simulation exercise.
Figure E15. Emo-Reframe: Generated by Care-Pilot using real data and domain knowledge.
[…]
Screenshot of the reframe pane in the simulation exercise.
E.2.2. Post-task Measurement and Pre-task Measurement